Determine the sum for the geometric series 1/2 - 1/3 + 2/9 - 4/27 +...
Answer by fractalier(2101)

(I fixed the 4th term...there was a typo)
The formula for finding the sum of an infinite geometric series is
S = a1 / (1-r)
Here a1 = 1/2 and r = -2/3 so that
S = (1/2) / (1 + 2/3)
S = (1/2) / (5/3)
S = 3/10